OpenStudy (anonymous):

find the average value of f(x) on the interval [-1, 4]

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

What is f(-1) ?

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

what is f(4) ?

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

So, \[\Large \text{Avg Rate of Change} = \frac{f(a)-f(b)}{a-b}\] \[\Large \text{Avg Rate of Change} = \frac{f(4)-f(-1)}{4-(-1)}\] \[\Large \text{Avg Rate of Change} = \frac{f(4)-f(-1)}{4+1}\] \[\Large \text{Avg Rate of Change} = \frac{f(4)-f(-1)}{5}\] \[\Large \text{Avg Rate of Change} = \frac{-1-0}{5}\] \[\Large \text{Avg Rate of Change} = -\frac{1}{5}\]